# Citadel: Context Similarity Based Deep Learning Framework Bug Finding

### v4

@@ 4.1. Preparation: Bug cases Collection
Fig. 3 shows a real report from the TensorFlow repository, consisting of the title, status, bug description, test code, and a timeline. To extract the buggy code, the sampler first judges whether one issue includes test code ( 1) and discards those lacking test code. Then, it checks issue labels ( 2) and discards reports that are not marked as bugs, crashes, etc. by developers. These issues usually report non-bug problems (e.g., documentation typos) and are not assigned labels by developers. During this process, the sampler also discards reports related to specific hardware (e.g., M1 chips (tor, 2023d)). Limited by the experiment environment, we cannot reproduce these bugs. Finally, the sampler examines the issue timeline ( 3) and discards issues with fewer than 3 comments, which are usually reports of issues that developers do not care about or intend to work on (tor, 2023b). It also discards closed issues that lack associated commits or pull requests. Such issues often arise from users’ misconceptions of expected behaviors and are promptly addressed by developers (tor, 2023c). After extracting buggy code, the sampler matches the most frequently mentioned API function from the ‘Title’ and ‘Bug Description’ (as shown in Fig. 3) as a problematic API candidate and verifies it in the corresponding test code. If the candidate does not appear in the code, the second most mentioned API is selected, and so on. Using such method, the problematic API in Fig. 3 can be correctly identified as tf.broadcast. Detailed implementation of the sampler is in our repository (our, 2023).
@@ 4.2. Context Similarity
Context similarity calculates the similarity between the runtime context of APIs, and the similar contexts intuitively show that the functionality of APIs would be similar. We observe that on some DL framework APIs (e.g., Conv1d, Conv2d, and Conv3d in PyTorch), although there are differences between their inputs or argument (e.g., different dimensions), they have similar functionalities and implementations, which can be reflected by context similarity. Many issues and patches (tor, 2022d, b, 2021) in GitHub further reveal that API functions that have similar functionalities are prone to have similar bugs due to one erroneous implementation of an underlying function. Moreover, the prior work (Wang et al., 2022) has proposed that the functional similarity between API functions can guide the construction of equivalence rules in testing. Therefore, we define and measure the context similarity between API functions to find API functions that have similar functionalities and leverage bugs on one API function to effectively identify potential bugs on its analogous API functions.
@@ 4.3. Dynamic Profiler
The most important context information of APIs is the source code functions they call during execution (i.e., the call stack), which intuitively demonstrates the underlying implementation of the API functions. To effectively collect such context information, the dynamic profiler executes unit test cases for API functions and record call stacks. The unit test cases consist of the test cases collected from DL framework repositories and the test cases generated by existing test case generation tools (Xie et al., 2022; Deng et al., 2022). These cases are intended to examine the expected behaviors of APIs during runtime and explore the analogous behaviors and edge cases. Fig. 4 illustrates how Citadel extract context information and matches Conv2d and LazyConvTranspose2d in the Fig. 2 as analogous APIs. The solid box in Fig. 4 illustrates part of the API call stacks collected by the profiler. During execution, both APIs call the source code function aten::convolution for performing convolution operations and cudaLaunchKernel related to GPU services. In addition, Conv2d calls the source code function aten::conv2d related to the optimization of 2D convolution, while LazyConvTranspose2d calls aten::conv_transpose2d related to transpose convolution.
@@ 4.4. Static Analyzer
Existing research (Assi et al., 2024) has revealed that a significant number of code blocks and functions with similar implementations exist within the source code of DL frameworks. Source code functions exhibiting similar functionality and implementation patterns may have similar errors, leading to bugs in the DL APIs that invoke them. The static analyzer is designed to identify and cluster these analogous functions within the source code. It then incorporates them as part of the context information provided to the API matcher, facilitating the measurement and matching of APIs with similar execution contexts. Inspired by the prior work (Alkhalaf et al., 2014; Su et al., 2016; McMillan et al., 2012; Novak et al., 2019), the static analyzer determines whether two functions, $F_{1}$ and $F_{2}$, have similar implementations by evaluating two key aspects: input and output arguments similarity $Sim_{io}$ and callees similarity $Sim_{call}$.
@@
Fig. 4. Matching APIs Pairs with Context Information
@@
otherwise, they are considered dissimilar. The dashed box in Fig. 4 shows how the static analyzer calculates the similarity between two source code functions aten::conv2d and aten::conv_transpose2d. Red marks input and output arguments and blue highlights callees. Given that these two functions share identical input and output arguments (e.g., const Tensor& weight) and call the same functions (e.g., at::convolution), both $Sim_{io}$ and $Sim_{call}$ exceed the thresholds. Therefore, aten::conv2d and aten::conv_transpose2d are classified into one group of analogous source code functions.
@@ 4.5. API matcher
The API matcher first receives context information from the dynamic profiler and static analyzer (i.e., analogous function groups and API call stacks) and matches analogous API pairs based on context similarity and signature similarity (Deng et al., 2022). Subsequently, it checks the arguments of analogous API functions and discards the API pairs with unsolvable argument mismatches which renders the buggy code unusable in test case generation. For instance, the API functions Conv2d and LPPool2d each have required arguments that are not included in the other, making Citadel unable to generate a test case for one based on the code of the other, as shown in Fig. 5 ( 4). Consequently, this API pair is considered to have encountered an argument mismatch and is discarded. Note that, matching and filtering API pairs in API matcher is a one-time process. The matched analogous API pairs can be saved and reused in the test case generator.
@@
where $f^{\prime}_{i}$ indicates the source code functions in the groups identified in the static analyzer and $f^{S}_{j}$ represents other source code functions in the call stack that do not belong to any of the groups. Similarly, we denote the execution context of another API function $A_{T}$ as $CTX_{T}$. When calculating the $Sim_{CTX}(A_{S},A_{T})$, if both APIs call source code functions from the same group, these functions will be treated as one function because of their similar functionality and implementations. The context similarity between $A_{S}$ and $A_{T}$ can be calculated by the Jaccard similarity coefficient:
@@
Fig. 4 illustrates the process of matching Conv2d and LazyConvTranspose2d mentioned in our motivaiton example via context information. Based on the records of the dynamic profiler (depicted in the solid box), both APIs call the source code functions aten::convolution and cudaLaunchKernel. For source code functions aten::conv2d and aten::conv_transpose2d which are separately called by Conv2d and LazyConvTranspose2d, they have been divided into one analogous function group by the static analyzer and are considered as the same function when calculating $Sim_{CTX}$. Since both APIs call almost the same source code functions, their Jaccard similarity coefficients are greater than $\beta$, and we match them as a context-similar API pair. In addition to context-similar API pairs, we supplement analogous API pairs matched by the signature similarity calculated by DeepREL (Deng et al., 2022), which are publicly available. Based on their experiment results, we select the top 20 API functions with the highest signature similarity to one API as its analogous API functions.
@@ 4.6. Test Case Generator
Dimension Difference. As mentioned previously, the DL framework provides a series of APIs for inputs of varying dimensions (e.g, Conv2d and Conv3d), typically sharing similar implementations and susceptibility to similar bugs (tor, 2022d). However, the existing methods encounter challenges in constructing test cases for these API functions due to the different dimensions of their arguments (Deng et al., 2022). To resolve the dimension difference, the test case generator first obtains the API signatures through open-source libraries (e.g., the ‘inspect’ library in Python) and identifies the dimension-related arguments from signatures. It then dynamically adjusts the test code by increasing or decreasing the dimensions of argument values based on the dimension information of the API signatures of the source and target APIs and generates test cases for those API functions. Fig. 5 shows an example of resolving dimension difference and generating available test cases for Conv3d ( 3). The API signature shows that the ‘input’ of Conv3d and Conv2d is dimension-related, and the ‘input’ of Conv2d is a 4-dimensional tuple, and the ‘input’ of Conv3d should be a 5-dimensional tuple. The generator recognizes such a dimension difference and expands the 4-dimension tuple received by Conv2d to a 5-dimension tuple to adjust the input dimensions (marked in green) and generates executable test cases for Conv3d.
@@ 4.7. Test Case Evaluator
$\bullet$ Performance bug arises from an underlying implementation or optimization error, leading to an unexpectedly high overhead on APIs. To detect performance bugs, Citadel calculates the expected overhead by leveraging the bug report of $A_{S}$ and records the actual runtime overhead. If $C_{T}$ exhibits unexpected overhead identical to that described in the bug report of the problematic API $A_{S}$ (e.g., a time overhead greater than expected), $A_{T}$ is considered to have a performance bug. Take the case in Fig. 2 as an example, the bug report of Conv2d indicates that the performance bug causes the group convolution to exhibit a greater time cost than implementing these convolutional layers individually, which is confirmed by the developers. Leveraging such an anomalous behavior as the test oracle, Citadel reveals that LazyConvTranspose2d also experiences a higher time overhead under the group setting than implementing LazyConvTranspose2d layers individually, thereby identifying the performance bug. Please note that this does not imply that LazyConvTranspose2d incurs the same cost as Conv2d in the original bug report, but rather that both achieve costs exceeding the cost of individual execution which is the expected overhead.
@@ 5.1. Setup
$\bullet$ Average time to detect bugs. Following the prior work (Cheu et al., 2003; Zhao et al., 2021), we use the metric of average time to detect bugs to compare the bug detection efficiency of each method. Specifically, we record the time taken by each method to conduct a complete test and subsequently divide it by the number of detected bugs in the execution. Since the manual cost of verifying which thousands of cases generated by baselines are real bugs is too high, for the baseline method, we directly use the total number of bugs reported in their papers. Note that the fuzzy-based baselines can explore more behaviors and trigger bugs through multiple executions. We actually calculate the upper bound of baselines’ performance in this metric, which is the time of a complete execution divided by the number of all bugs they report. In contrast, Citadel exploits existing bug reports to discover new bugs without fuzzy generation, and all bugs reported by Citadel are discovered in one execution.
@@
Collected Issues and Context Information. In the experiment, we separately collect and verify 258 and 288 valid bug cases from the latest 30,000 issues of PyTorch and the latest 20,000 issues of TensorFlow, as one input to Citadel. This process takes approximately 3 weeks per participant. Among them, the problematic APIs in 172 cases (i.e., 104 PyTorch cases and 68 TensorFlow cases) have matched analogous APIs and are used to generate new test cases in the test case generator. These bug cases are publicly available in our repository (our, 2023). In addition, the dynamic profiler constructs test cases for 999 PyTorch APIs and 2,076 TensorFlow APIs and records their call stacks. Based on the experimental results of prior work (Su et al., 2016; Misu and Sakib, 2017), we set $\alpha_{1}=0.8$ as the threshold of input and output arguments similarity. To strictly judge analogous source code functions and reduce the impact of false positives on subsequent API matching, we set $\alpha_{2}=0.8$ as the threshold for callees similarity. If both the arguments similarity and callees similarity of two functions exceed the thresholds, the two source code functions are judged as analogous functions. The static analyzer separately selects 944 and 7,028 functions in PyTorch and TensorFlow source code that share the similarity with at least one other function from the source code and divides them into 2,467 groups. The call stacks and analogous function groups are input into the API matcher as context information.
@@ 5.2. Effectiveness in Detecting Bugs
Experiment Design: To evaluate the effectiveness of Citadel in detecting real-world bugs, we conduct experiments on the PyTorch and TensorFlow frameworks and report all detected API bugs to developers for confirmation. Our experiment counts the number of API bugs detected by Citadel, that is, when calling an API with certain inputs triggers one bug, Citadel will consider that an API bug is detected. In addition, aligned to prior work (Xie et al., 2022), different inputs triggering the same unexpected behavior on one API will only be counted once. During this process, we record the states of reports, such as confirmed or rejected, and the bug types. Furthermore, following the setting of previous work (Deng et al., 2022; Wei et al., 2022), we collect the number of verified pairs and covered APIs in the API matcher of Citadel. Note that Citadel uses the test code in existing bug reports to generate new test cases, and the number of APIs tested in the experiment is related to the collected bug cases. We also report the number of tested APIs using the collected 172 bug cases. How to collect more bug cases to fully utilize the matched API pairs and test more APIs will be a future research direction.
Results: Table 1 summarizes the three types of bugs detected by Citadel, namely ‘Stat.’, ‘Val.’, and ‘Perf.’ (i.e., status, value, and performance bugs). The first column of Table 1 displays the DL framework and the following columns indicate the number of all detected bugs (‘#Total’), bugs that developers do not plan to work on (‘#Rejected’), bugs that are duplicates of existing bug reports (‘#Duplicated’), new bugs that have not been reported (‘#New’), and new bugs that have been confirmed (‘#Confirmed’). In addition, Table 2 presents a comparison between Citadel and baselines in the three metrics in § 5.1. The first column shows four test approaches in comparison and the second column displays the DL framework. The columns ‘#API’ and ‘#Pairs’ show the number of covered APIs and matched API pairs in each approach, corresponding to the ‘Number of covered APIs’ in § 5.1. In addition, Fig. 6 uses Venn diagrams to present the comparison of the number of covered APIs and matched API pairs between Citadel and the baselines. Since we could not find specifics on the APIs and pairs covered by DeepREL in its repository, we use the data collected from its full execution.
@@
Table 2. Comparison of Citadel and Baselines Result
@@
Citadel PyTorch 1,436 8,079 82 196 41.84 4.83
TensorFlow 5,380 28.268 61 208 29.33 6.61
Total 6,816 36,347 143 404 35.40 5.70
@@
Analysis: The results in Table 1 illustrate the effectiveness of Citadel in detecting various types of real-world bugs. Citadel generates test cases based on a total of 172 real bugs collected from GitHub repositories and successfully detects 151 API bugs on PyTorch and TensorFlow, out of which only 3 are rejected by developers. These cases show the same anomalous behaviors (e.g., NaN) as the source problematic API, but developers have no plans to fix them. The following ‘Rejected Case’ provides a detailed analysis of a rejected case. Of the remaining 148 bugs, 24 are duplicates of existing bug reports and the remaining 124 are new bugs, 92 of which have been confirmed by developers. Excluding rejected cases, Citadel detected 101 status bugs, 34 value bugs, and 13 performance bugs, demonstrating its effectiveness in detecting different types of DL framework bugs. Compared to other types of bugs, the number of detected performance bugs is relatively small. Our manual analysis of the PyTorch and TensorFlow repositories reveals that the reported performance bugs are infrequent. Take the PyTorch framework as an example, only approximately 300 issues are labeled as ‘performance’ out of over 10,000 open issues. Moreover, the limited number of issues with reproducible code poses a challenge for Citadel to gather a significant amount of code related to performance bugs in DL framework repositories. Finally, 8/172 real bugs that Citadel collects to generate test cases for the matched API pairs are related to performance. Based on these collected performance bugs, Citadel detects 13 new performance bugs, and 11 of them have been confirmed.
@@
The API matcher of Citadel covers a total of 1,436 PyTorch APIs and 5,380 TensorFlow APIs, which is 365 more PyTorch APIs and 3,478 more TensorFlow APIs than DeepREL’s ‘API_Match_Verifier’ and significantly more than 498 PyTorch APIs and 911 TensorFlow APIs covered by DocTer’s constraints. Fig. 6 show the comparison of API and API pairs covered by Citadel and baselines, respectively. In addition, the experiments on the collected bug cases cover 529 PyTorch APIs and 675 TensorFlow APIs, of which 221 and 338 are not covered by the baselines. Citadel identifies one PyTorch bug and eight TensorFlow bugs on these APIs. Moreover, these test cases cover 797 and 1,387 analogous API pairs on PyTorch and TensorFlow, of which 574 and 1,154 are not covered by the baselines. The newly covered API pairs enable Citadel to successfully detect 49 API bugs on PyTorch and TensorFlow, comprising 41 status bugs, 3 value bugs, and 5 performance bugs, which demonstrates the effectiveness of the newly covered APIs and API pairs of Citadel in detecting bugs. For example, Citadel newly matches the API pair of Conv2d and LazyConvTranspose2d via the context similarity and identifies the performance bug in Fig. 2. The API pair of ReLU6 and HardTanh in the following case study is also newly covered in Citadel.

### v5

@@ 4.1. Preparation: Bug cases Collection
Fig. 3 shows a real report from the TensorFlow repository, consisting of the title, status, bug description, test code, and a timeline. To extract the buggy code, the sampler first judges whether one issue includes test code ( 1) and discards those lacking test code. Then, it checks issue labels ( 2) and discards reports that are not marked as bugs, crashes, etc. by developers. These issues usually report non-bug problems (e.g., documentation typos) and are not assigned labels by developers. During this process, the sampler also discards reports related to specific hardware (e.g., M1 chips (tor, 2023d)). Limited by the experiment environment, we cannot reproduce these bugs. Finally, the sampler examines the issue timeline ( 3) and discards issues with fewer than 3 comments, which are usually reports of issues that developers do not care about or intend to work on (tor, 2023b). It also discards closed issues that lack associated commits or pull requests. Such issues often arise from users’ misconceptions of expected behaviors and are promptly addressed by developers (tor, 2023c). After extracting buggy code, the sampler matches the most frequently mentioned API function from the ‘Title’ and ‘Bug Description’ (as shown in Fig. 3) as a problematic API candidate and verifies it in the corresponding test code. If the candidate does not appear in the code, the second most mentioned API is selected, and so on. With such a method, the problematic API in Fig. 3 can be correctly identified as tf.broadcast. Detailed implementation of the sampler is in our repository (our, 2023).
@@
We acknowledge that, similar to prior approaches, CITADEL also requires a certain amount of manual effort, mainly to verify and ensure the effectiveness of the test cases extracted from bug reports, as described in Section 4.1..
The manual effort at this preparation stage is mainly to verify and ensure the effectiveness of the test cases extracted from bug reports, which will be used in the following experiments (§ 5.2). Note that in some application scenarios, once the effectiveness of the test cases is ensured, the associated manual effort can be significantly reduced or even eliminated. For example, when Citadel is used in conjunction with other fuzzing tools, it can directly leverage the test cases that are generated by other tools and have triggered individual API bugs as input. Citadel can automatically construct new test cases for analogous APIs without requiring additional verification, thereby enabling effective and efficient bug detection.
@@ 4.2. Context Similarity
Context similarity calculates the similarity between the runtime context of APIs, and the similar contexts intuitively show that the functionality of APIs would be similar. We observe that on some DL framework APIs (e.g., Conv1d, Conv2d, and Conv3d in PyTorch), although there are differences between their inputs or arguments (e.g., different dimensions), they have similar functionalities and implementations, which can be reflected by context similarity. Many issues and patches (tor, 2022d, b, 2021) in GitHub further reveal that API functions that have similar functionalities are prone to have similar bugs due to one erroneous implementation of an underlying function. Moreover, the prior work (Wang et al., 2022) has proposed that the functional similarity between API functions can guide the construction of equivalence rules in testing. Therefore, we define and measure the context similarity between API functions to find API functions that have similar functionalities and leverage bugs on one API function to effectively identify potential bugs on its analogous API functions.
@@ 4.3. Dynamic Profiler
The most important context information of APIs is the source code functions they call during execution (i.e., the call stack), which intuitively demonstrates the underlying implementation of the API functions. To effectively collect such context information, the dynamic profiler executes unit test cases for API functions and records call stacks. The unit test cases consist of the test cases collected from DL framework repositories and the test cases generated by existing test case generation tools (Xie et al., 2022; Deng et al., 2022). These cases are intended to examine the expected behaviors of APIs during runtime and explore the analogous behaviors and edge cases. Fig. 4 illustrates how Citadel extract context information and matches Conv2d and LazyConvTranspose2d in the Fig. 2 as analogous APIs. The solid box in Fig. 4 illustrates part of the API call stacks collected by the profiler. During execution, both APIs call the source code function aten::convolution for performing convolution operations and cudaLaunchKernel related to GPU services. In addition, Conv2d calls the source code function aten::conv2d related to the optimization of 2D convolution, while LazyConvTranspose2d calls aten::conv_transpose2d related to transpose convolution.
@@ 4.4. Static Analyzer
Existing research (Assi et al., 2025) has revealed that a significant number of code blocks and functions with similar implementations exist within the source code of DL frameworks. Source code functions exhibiting similar functionality and implementation patterns may have similar errors, leading to bugs in the DL APIs that invoke them. The static analyzer is designed to identify and cluster these analogous functions within the source code. It then incorporates them as part of the context information provided to the API matcher, facilitating the measurement and matching of APIs with similar execution contexts. Inspired by the prior work (Alkhalaf et al., 2014; Su et al., 2016; McMillan et al., 2012; Novak et al., 2019), the static analyzer determines whether two functions, $F_{1}$ and $F_{2}$, have similar implementations by evaluating two key aspects: input and output arguments similarity $Sim_{io}$ and callees similarity $Sim_{call}$.
@@
Fig. 4. A Demo Case of Matching APIs Pairs with Context Information
@@
otherwise, they are considered dissimilar. The dashed box in Fig. 4 shows how the static analyzer calculates the similarity between two source code functions aten::conv2d and aten::conv_transpose2d. Red marks input and output arguments, and blue highlights callees. Given that these two functions share identical input and output arguments (e.g., const Tensor& weight) and call the same functions (e.g., at::convolution), both $Sim_{io}$ and $Sim_{call}$ exceed the thresholds. Therefore, aten::conv2d and aten::conv_transpose2d are classified into one group of analogous source code functions.
@@ 4.5. API matcher
The API matcher first receives context information from the dynamic profiler and static analyzer (i.e., analogous function groups and API call stacks) and matches analogous API pairs based on context similarity and signature similarity (Deng et al., 2022). Subsequently, it checks the arguments of analogous API functions and discards the API pairs with unsolvable argument mismatches, which renders the buggy code unusable in test case generation. For instance, the API functions Conv2d and LPPool2d each have required arguments that are not included in the other, making Citadel unable to generate a test case for one based on the code of the other, as shown in Fig. 5 ( 4). Consequently, this API pair is considered to have encountered an argument mismatch and is discarded. Note that matching and filtering API pairs in API matcher is a one-time process. The matched analogous API pairs can be saved and reused in the test case generator.
@@
where $f^{\prime}_{i}$ indicates the source code functions in the groups identified in the static analyzer and $f^{S}_{j}$ represents other source code functions in the call stack that do not belong to any of the groups. Similarly, we denote the execution context of another API function $A_{T}$ as $CTX_{T}$. Note that Citadel does not merely compare whether the two API call stacks are the same. When calculating the $Sim_{CTX}(A_{S},A_{T})$, if both APIs call source code functions from the same group, these functions will be treated as one function because of their similar functionality and implementations. Such a design enables Citadel to match APIs with similar underlying implementations but different call stacks. The context similarity between $A_{S}$ and $A_{T}$ can be calculated by the Jaccard similarity coefficient:
@@
Fig. 4 illustrates a simplified process of matching Conv2d and LazyConvTranspose2d mentioned in our motivation example via context information. Based on the records of the dynamic profiler (depicted in the solid box), both APIs call the source code functions aten::convolution and cudaLaunchKernel. For source code functions aten::conv2d and aten::conv_transpose2d, which are separately called by Conv2d and LazyConvTranspose2d, they have been divided into one analogous function group by the static analyzer and are considered as the same function when calculating $Sim_{CTX}$. Note that there are multiple differences in the call stacks of Conv2d and LazyConvTranspose2d, aten::conv2d and aten::conv_transpose2d are two demo cases. Since both APIs call almost the same source code functions, their Jaccard similarity coefficients are greater than $\beta$, and we match them as a context-similar API pair. In addition to context-similar API pairs, we supplement analogous API pairs matched by the signature similarity calculated by DeepREL (Deng et al., 2022), which are publicly available. Based on their experiment results, we select the top 20 API functions with the highest signature similarity to one API as its analogous API functions.
@@ 4.6. Test Case Generator
Dimension Difference. As mentioned previously, the DL framework provides a series of APIs for inputs of varying dimensions (e.g, Conv2d and Conv3d), typically sharing similar implementations and susceptibility to similar bugs (tor, 2022d). However, the existing methods encounter challenges in constructing test cases for these API functions due to the different dimensions of their arguments (Deng et al., 2022). To resolve the dimension difference, the test case generator first obtains the API signatures through open-source libraries (e.g., the ‘inspect’ library in Python) and identifies the dimension-related arguments from signatures. It then dynamically adjusts the test code by increasing or decreasing the dimensions of argument values based on the dimension information of the API signatures of the source and target APIs, and generates test cases for those API functions. Fig. 5 shows an example of resolving dimension difference and generating available test cases for Conv3d ( 3). The API signature shows that the ‘input’ of Conv3d and Conv2d is dimension-related, and the ‘input’ of Conv2d is a 4-dimensional tuple, and the ‘input’ of Conv3d should be a 5-dimensional tuple. The generator recognizes such a dimension difference and expands the 4-dimension tuple received by Conv2d to a 5-dimension tuple to adjust the input dimensions (marked in green) and generates executable test cases for Conv3d.
@@ 4.7. Test Case Evaluator
$\bullet$ Performance bug arises from an underlying implementation or optimization error, leading to an unexpectedly high overhead on APIs. To detect performance bugs, Citadel calculates the expected overhead by leveraging the bug report of $A_{S}$ and records the actual runtime overhead. If $C_{T}$ exhibits unexpected overhead identical to that described in the bug report of the problematic API $A_{S}$ (e.g., a time overhead greater than expected), $A_{T}$ is considered to have a performance bug. Take the case in Fig. 2 as an example, the bug report of Conv2d indicates that the performance bug causes the group convolution to exhibit a greater time cost than implementing these convolutional layers individually, which is confirmed by the developers. Leveraging such an anomalous behavior as the test oracle, Citadel reveals that LazyConvTranspose2d also experiences a higher time overhead under the group setting than implementing LazyConvTranspose2d layers individually, thereby identifying the performance bug. Please note that this does not imply that LazyConvTranspose2d incurs the same cost as Conv2d in the original bug report, but rather that both achieve costs exceeding the cost of individual execution (i.e., the expected overhead).
@@ 5.1. Setup
$\bullet$ Average time to detect bugs. Following the prior work (Cheu et al., 2003; Zhao et al., 2021), we use the metric of average time to detect bugs to compare the bug detection efficiency of each method. Specifically, we record the time taken by each method to conduct a complete test run and subsequently divide it by the number of detected bugs in the execution. Note that this metric explicitly excludes any preprocessing time, such as API matching in CITADEL and DeepREL or constraint construction in DocTer, since these steps often involve manual effort or produce static results that can be reused in multiple tests. Since the manual cost of verifying which thousands of cases generated by baselines are real bugs is too high, for the baseline method, we directly use the total number of bugs reported in their papers. Note that the fuzzy-based baselines can explore more behaviors and trigger bugs through multiple executions. We actually calculate the upper bound of baselines’ performance in this metric, which is the time of a complete execution divided by the total number of bugs they report. In contrast, Citadel exploits existing bug reports to discover API bugs without fuzzy generation, and all bugs reported by Citadel are discovered in one execution.
@@
Collected Issues and Context Information. In the experiment, we separately collect and verify 258 and 288 valid bug cases from the latest 30,000 issues of PyTorch and the latest 20,000 issues of TensorFlow, as one input to Citadel. This process takes approximately 3 weeks per participant. Among them, the problematic APIs in 172 cases (i.e., 104 PyTorch cases and 68 TensorFlow cases) have matched analogous APIs and are used to generate new test cases in the test case generator. These bug cases are publicly available in our repository (our, 2023). In addition, the dynamic profiler constructs test cases for 999 PyTorch APIs and 2,076 TensorFlow APIs and records their call stacks. Based on the experimental results of prior work (Su et al., 2016; Misu and Sakib, 2017), we set $\alpha_{1}=0.8$ as the threshold of input and output arguments similarity. To strictly judge analogous source code functions and reduce the impact of false positives on subsequent API matching, we set $\alpha_{2}=0.8$ as the threshold for callees similarity. If both the arguments similarity and callees similarity of two functions exceed the thresholds, the two source code functions are judged as analogous functions. The static analyzer separately selects 944 and 7,028 functions in PyTorch and TensorFlow source code that share similarity with at least one other function from the source code and divides them into 2,467 groups. The call stacks and analogous function groups are input into the API matcher as context information.
@@ 5.2. Effectiveness in Detecting Bugs
Experiment Design: To evaluate the effectiveness of Citadel in detecting real-world bugs, we conduct experiments on the PyTorch and TensorFlow frameworks and report all detected API bugs to developers for confirmation. Our experiment counts the number of API bugs detected by Citadel, that is, when calling an API with certain inputs triggers one bug, Citadel will consider that an API bug is detected. In addition, aligned to prior work (Xie et al., 2022), different inputs triggering the same unexpected behavior on one API will only be counted once. During this process, we record the states of reports, such as confirmed or rejected, and the bug types. Furthermore, following the setting of previous work (Deng et al., 2022; Wei et al., 2022), we collect the number of verified pairs and covered APIs in the API matcher of Citadel. Note that Citadel uses the test code in existing bug reports to generate new test cases, and the number of APIs tested in the experiment is related to the collected bug cases. We also report the number of tested APIs using the collected 172 bug cases. How to collect more bug cases to fully utilize the matched API pairs and test more APIs will be a future direction.
Results: Table 1 summarizes the three types of bugs detected by Citadel, namely ‘Stat.’, ‘Val.’, and ‘Perf.’ (i.e., status, value, and performance bugs). Following the prior work (Xie et al., 2022; Deng et al., 2022, 2023), the number of bugs reported here is the number of abnormal behaviors in different APIs (i.e., API bugs), rather than the number of independent implementation defects. The first column of Table 1 displays the DL framework and the following columns indicate the number of all detected API bugs (‘#Total’), API bugs that developers do not plan to work on (‘#Rejected’), API bugs that have been reported in existing reports (‘#Duplicated’), new API bugs that have not been reported (‘#New’), new API bugs that have been confirmed (‘#Confirmed’). In addition, Table 2 presents a comparison between Citadel and baselines in the three metrics in § 5.1. The first column shows four test approaches in comparison and the second column displays the DL framework. The columns ‘#API’ and ‘#Pairs’ show the number of covered APIs and matched API pairs in each approach, corresponding to the ‘Number of covered APIs’ in § 5.1. In addition, Fig. 6 uses Venn diagrams to present the comparison of the number of covered APIs and matched API pairs between Citadel and the baselines. Since we could not find specifics on the APIs and pairs covered by DeepREL in its repository, we use the data collected from its full execution.
@@
Table 2. Comparison of Citadel and Baselines Result (Brackets mark the APIs and API Pairs Covered by the Collected Bug Cases)
@@
Citadel PyTorch 1,436 (529) 8,079(797) 82 196 41.84 4.83
TensorFlow 5,380 (675) 28,268 (1,387) 61 208 29.33 6.61
Total 6,816 (1,204) 36,347 (2,184) 143 404 35.40 5.70
@@
Analysis: The results in Table 1 illustrate the effectiveness of Citadel in detecting various types of real-world bugs. Citadel generates test cases based on a total of 172 real bugs collected from GitHub repositories and successfully detects 151 API bugs on PyTorch and TensorFlow, out of which only 3 are rejected by developers. These cases show the same anomalous behaviors (e.g., NaN) as the source problematic API, but developers have no plans to fix them. The following ‘Rejected Case’ provides a detailed analysis of a rejected case. Of the remaining 148 bugs, 24 are duplicates of existing bug reports and the remaining 124 are unreported API bugs, 92 of which have been confirmed by developers. Excluding rejected cases, Citadel detected 101 status bugs, 34 value bugs, and 13 performance bugs, demonstrating its effectiveness in detecting different types of DL framework bugs. Furthermore, we analyze the patches provided by developers in response to our bug reports and count the number of API bugs that share the same patch as the source API bug. In PyTorch, eight of our issue reports receive official patches (including twelve API bugs), while the patches in six reports indicate that seven newly reported API bugs share the same underlying implementation errors with their corresponding source bugs. In addition, due to the inactivity of the TensorFlow community, none of our bug reports have received patches, making this analysis impossible. This finding highlights a characteristic of our context similarity-based method in Citadel, i.e., it could identify multiple API bugs caused by the same underlying implementation error. Moreover, compared to other types of bugs, the number of detected performance bugs is relatively small. Our manual analysis of the PyTorch and TensorFlow repositories reveals that the reported performance bugs are infrequent. Take the PyTorch framework as an example, only approximately 300 issues are labeled as ‘performance’ out of over 10,000 open issues. Moreover, the limited number of issues with reproducible code poses a challenge for Citadel to gather a significant amount of code related to performance bugs in DL framework repositories. Finally, 8/172 real bugs that Citadel collects to generate test cases for the matched API pairs are related to performance. Based on these collected performance bugs, Citadel detects 13 new performance bugs, and 11 of them have been confirmed.
@@
The API matcher of Citadel covers a total of 1,436 PyTorch APIs and 5,380 TensorFlow APIs, which is 365 more PyTorch APIs and 3,478 more TensorFlow APIs than DeepREL’s ‘API_Match_Verifier’ and significantly more than 498 PyTorch APIs and 911 TensorFlow APIs covered by DocTer’s constraints. Fig. 6 show the comparison of API and API pairs covered by Citadel and baselines, respectively. In addition, the experiments based on the collected bug cases cover 529 PyTorch APIs and 675 TensorFlow APIs, among which 221 and 338, respectively, are not covered by the baseline methods. Citadel successfully identifies one PyTorch bug and eight TensorFlow bugs on these previously uncovered APIs. Furthermore, the corresponding test cases cover 797 and 1,387 analogous API pairs on PyTorch and TensorFlow, of which 574 and 1,154 are not covered by the baselines. Note that, due to the limited number of collected bug cases, the experiment currently only uses a small subset of matched analogous APIs and API pairs. Exploring techniques to automatically annotate and efficiently collect a broader set of bug cases to further enhance the detection effect of Citadel will be our future research direction (§ 7). The newly covered API pairs enable Citadel to successfully detect 49 API bugs on PyTorch and TensorFlow, comprising 41 status bugs, 3 value bugs, and 5 performance bugs, which demonstrates the effectiveness of the newly covered APIs and API pairs of Citadel in detecting bugs. For example, Citadel newly matches the API pair of Conv2d and LazyConvTranspose2d via the context similarity and identifies the performance bug in LazyConvTranspose2d (Fig. 2). The API pair of ReLU6 and HardTanh in the following case study is also newly covered in Citadel.